OpenStudy (anonymous):

Which of the following is a true statement? a. 8%=0.8=4/5 b. 2/5=0.4=40% c. 0.06=6/10=6% d. 3/8=.375=375%

OpenStudy (anonymous):

b.

OpenStudy (anonymous):

a. 0.8 *100 = 80% not 8 c. 6/10 = 0.6 not .06 d. 0.375 * 100 = 37.5% not 375%